OpenStudy (anonymous):

How would i simplify this algebraic expression (3ab^2) * (-2a^4b)

OpenStudy (jdoe0001):

\(\bf (3ab^2) \times (-2a^4b)\implies 3\times -2\times a^1\times a^4\times b^2\times b^4\\ \quad \\ \large -6a^{1+4}b^{2+4}\)